OpenStudy (anonymous):

Does the rule y=4.4^x represent a linear or an exponential function? Explain

OpenStudy (amistre64):

i would assume if it has an exponent in it; that it might possibly be exponential. I have read some other definitions of an exponential function, but i think this is more in line with the question tho.

OpenStudy (anonymous):

I would say it is an exponential functions. Since it is raised to the power of x, each time the exponent will be greater. y(1)= 4.4; y(2)=4.4^2; y(3)=4.4^3 and so on. A linear function would be something times x. or any variable. But it would literally be just a line in the form of A*x+b where A is the coefficient and b just any number.